OpenStudy (anonymous):

For a class pizza party, Esther bought twelvesmall bottles of juice and 9 large bottles of water. Let j represent the cost of juice and w represent the cost of water. Write an expression to determine the total cost for the drinks. Then, calculate the cost if juice costs one dollar and water costs two dollars.

OpenStudy (anonymous):

Hello, Twilaswift. Lovely name. (: To begin, my equation will be j*12 + w*9 = total. Why? Simply because J = the juice's cost, we're buying 12 bottles of juice (that's a lot of juice.) W = Water, and we're buying 9 large bottles of water. So if we want the total then... J*12 + W*9 = Total or.. X. You can also write this as J12 + W9 = X Now, if we know that juice is $1 a piece, and water is $2 a piece then it get pretty simple here. We replace J in J*12 with 1. 1*12 = 12. So $12 for the juice. As for the water, it is $2 per piece, so W replaced with 2 would be... 2*9, which equals 18. $18. (: So we just add up our two final figures of $12 and $18. $12 + $18 = $30. :D Did that help? :D